1. In a study of the formation of NOx air pollution, a chamber heated to 2200°C was filled with air (0.790 atm N2, 0.210 atm O2). What are the equilibrium partial pressures of N2, O2, and NO if Kp = 0.0510 for the following reaction:
N2(g) + O2 (g)⇌ 2NO (g)
1
Expert's answer
2020-04-13T02:25:36-0400

Initial:

0.790 atm N2, 0.210 atm O2, 0 atm NO.

Reacted:

x N2, x O2.

Equilibrium:

(0.790 - x) atm N2, (0.210 - x) atm O2, 2x atm NO

Kp = [NO]2/([N2]*[O2])

0.0510 = (2x)2 / ((0.790 - x)*(0.210 - x))

From the equation:

x = 0.040 atm - reacted

Equilibrium:

0.790 - x = 0.0790 - 0.040 = 0.750 atm N2

0.210 - x = 0.210 - 0.040 = 0.170 atm O2

2x = 0.040 * 2 = 0.080 atm NO